Karen’s father is 7 times as old as Karen is. One year ago he was 9 times as old as Karen was. Find Karen’s present age.
Answer by ikleyn(53762)

.

Let x = Karen's age.

Then father's age is 7x.


One year ago


    9*(x-1) = 7x-1.


Simplify and solve


    9x - 9 = 7x - 1

    9x - 7x = 9 - 1

    2x      = 8

     x      = 4


ANSWER.  Karen's present age is 4 years.

Solved.